# Stochastic weight matrix dynamics during learning and Dyson Brownian motion

### v1

@@ 3.2 Universal predictions
Importantly, from the general discussion in the previous section it follows that the higher the level of stochasticity (as indicated by $g_{i}^{2}$ and depending on learning rate and batch size), the wider the distributions are expected to be. We have hence repeated the exercise above for a range of learning rates, $0.01\leq\alpha\leq 0.1$, and mini-batch sizes, $4\leq|{\cal B}|\leq 64$. We will demonstrate below that the potentials $V_{i}(x_{i})$ have a global minimum at $x_{i}^{s}=\mu^{2}-\kappa_{i}$. The variance in the Coulomb gas $\sigma_{i}^{2}$ is then proportional to the ratio of the learning rate and batch size, see Eq. (2.18), and both the mean level spacing $\langle S\rangle$ and the width of the spectral density are expected to scale with $\sqrt{\alpha/|{\cal B}|}$ times a model and data dependent function of $\mu^{2}$ and $\kappa_{i}$.
@@
$\langle S\rangle,\sqrt{\pi}\sigma=a_{\rm fit}\sqrt{\frac{\alpha}{|{\cal B}|}\kappa_{i}^{2}\Omega_{i}}.$ (3.10)
A derivation of the non-universal factor $\kappa_{i}^{2}\Omega_{i}=(\mu^{2}-\kappa_{i})$, using an analysis of the variance of the gradient of the loss function, can be found in App. C. A linear dependence of $\langle S\rangle^{2}$ and $\sigma^{2}$ on $\alpha/|{\cal B}|$ is observed, consistent with the derivation in Sec. 2.
@@
3.3 Non-universal dynamics
Next we turn to non-universal, model-dependent features, which are contained in the drift $K_{i}(x_{i})$ and the potential $V_{i}(x_{i})$. In Refs. [21, 22] it was shown that in the instantaneous eigenbasis the singular values $\xi_{i}$ of $W$ are subject to the equation
$\frac{d}{dt}\xi_{i}=\sigma_{h}^{2}\left(\frac{1}{\kappa_{i}}-\frac{1}{\mu^{2}-\sigma_{h}^{2}\xi_{i}^{2}}\right)\xi_{i},$ (3.11)
@@ 3.3 Non-universal dynamics
$\frac{d}{d\tau}x_{i}=\left(\frac{1}{\kappa_{i}}-\frac{1}{\mu^{2}-x_{i}}\right)x_{i}.$ (3.12)
However, it should be clear from the previous discussion that this equation is not relevant for the evolution of the eigenvalues $x_{i}$ of $X$ in the realistic case, as it lacks both the Coulomb repulsion and the stochasticity. Instead, we should consider – see Eq. (A.2) –
$\frac{d}{d\tau}x_{i}=K_{i}(x_{i})+\sum_{j\neq i}\frac{g_{i}^{2}}{x_{i}-x_{j}}+\sqrt{2}g_{i}\eta_{i},$ (3.13)
@@
$K_{i}(x_{i})=\left(\frac{1}{\kappa_{i}}-\frac{1}{\mu^{2}-x_{i}}\right)x_{i}.$ (3.14)
@@
$\partial_{\tau}P(x,\tau)=\partial_{x}\left(g^{2}\partial_{x}-K(x)\right)P(x,\tau),\qquad\qquad K(x)=\left(\frac{1}{\kappa}-\frac{1}{\mu^{2}-x}\right)x.$ (3.15)
@@
$P_{s}(x)=\frac{1}{Z}e^{-V(x)/g^{2}}=\frac{1}{Z}\exp\left[\frac{1}{g^{2}}\left(\frac{x^{2}}{2\kappa}+x+\mu^{2}\log\left(\mu^{2}-x\right)\right)\right].$ (3.17)
@@
To analyse the properties of the time-dependent FPE, we cast the dynamics as a quantum-mechanical bound state problem [23]. We factor out the square root of the stationary distribution
$P(x,\tau)=\sqrt{P_{s}(x)}\psi(x,\tau),$ (3.19)
and analyse the time dependence of $\psi(x,\tau)$. Taking the time derivative and using Eq. (3.15) then yields
$\partial_{\tau}\psi(x,\tau)=\left(g^{2}\partial_{x}^{2}-\frac{1}{4g^{2}}\left[\partial_{x}V(x)\right]^{2}+\frac{1}{2}\left[\partial_{x}^{2}V(x)\right]\right)\psi(x,\tau)\equiv-2H_{\rm FP}\psi(x,\tau),$ (3.20)
@@
$L^{\dagger}=-g\partial_{x}+\frac{1}{2g}\partial_{x}V(x),\qquad\qquad L=+g\partial_{x}+\frac{1}{2g}\partial_{x}V(x),$ (3.22)
@@
$H_{\rm FP}\psi_{n}(x)=E_{n}\psi_{n}(x).$ (3.23)
The ground state, $\psi_{0}(x)$, with vanishing energy $E_{0}=0$, is determined by $L\psi_{0}(x)=0$, which yields $\psi_{0}(x)=\sqrt{P_{s}(x)}$, as it should be, see Eqs. (3.17) and (3.19).
@@
$\psi(x,\tau)=\psi_{0}(x)+\sum_{n>0}c_{n}\psi_{n}(x)e^{-2E_{n}\tau},$ (3.24)
where the coefficients $c_{n}$ are determined by the initial distribution $P(x,0)$, using relation (3.19).
@@
$H_{\rm FP}=-\frac{g^{2}}{2}\partial_{x}^{2}+U(x),\qquad\qquad U(x)=\frac{1}{g^{2}}\left[U_{0}(x)+g^{2}U_{1}(x)\right],$ (3.25)
@@
$U_{0}(x)=\frac{1}{8}\left[\partial_{x}V(x)\right]^{2},\qquad\qquad U_{1}(x)=-\frac{1}{4}\partial_{x}^{2}V(x),$ (3.26)
@@
$\partial_{x}U_{0}(x)\big{|}_{x=x_{s}}=0,\qquad\qquad\partial_{x}V(x)\big{|}_{x=x_{s}}=-K(x_{s})=0,$ (3.27)
@@
$\partial_{x}^{2}U_{0}(x)\big{|}_{x=x_{s}}=\frac{\Omega^{2}}{4},\qquad\qquad\Omega\equiv\partial_{x}^{2}V(x)\big{|}_{x=x_{s}}=\frac{\mu^{2}-\kappa}{\kappa^{2}}.$ (3.28)
@@
$H_{\rm FP}=-\frac{g^{2}}{2}\partial_{x}^{2}+\frac{1}{2g^{2}}\left(\frac{\Omega}{2}\right)^{2}\left(x-x_{s}\right)^{2}-\frac{1}{2}\left(\frac{\Omega}{2}\right),$ (3.29)
@@
$E_{n}=\frac{\Omega}{2}n,\qquad\quad n=0,1,2,\ldots,$ (3.30)
and convergence to the stationary limit is determined by the smallest nonzero eigenvalue, $2E_{1}=\Omega$, as observed before [22]. Also shown in Fig. 8 is the groundstate wave function, $\psi_{0}(x)=\sqrt{P_{s}(x)}$. We have a chosen a rather large value of $g^{2}$ here, such that it has a substantial width. Indeed, in the harmonic approximation, $g^{2}$ does not appear in the spectrum but controls the width of the wave function, and hence the distribution $P(x,t)$. Less stochasticity results in a narrower distribution centred around $x=x_{s}$, as one would expect.
@@
3.4 Teacher-student model
@@
$\mathbf{y}_{t}=V\mathbf{x},\qquad\qquad\mathbf{y}_{s}=W\mathbf{x},$ (3.31)
@@
${\cal L}(W)=\frac{1}{2}\left\langle||\mathbf{y}_{t}-\mathbf{y}_{s}||^{2}\right\rangle,$ (3.32)
where the brackets denote averaging over the input data. In the limit of an infinite amount of input data, $\langle x_{i}x_{j}\rangle=\delta_{ij}$, and the dynamics is deterministic. Some stochasticity can be introduced by considering batch updates. The amount of stochasticity is, however, not sufficient for the purpose of this discussion, since it is unavoidably suppressed when $W\to V$ as the training converges. We therefore introduce additional randomness by hand, in a manner that is consistent with the discussion in Sec. 2, see in particular Eq. (2.5). We consider
$W_{ij}\to W_{ij}^{\prime}=W_{ij}-\alpha\frac{\delta{\cal L}}{\delta W_{ij}}+\alpha\sqrt{A_{ij}}\eta_{ij}^{\cal B}.$ (3.33)
@@
$\langle\eta_{ij}^{\cal B}\eta_{kl}^{\cal B}\rangle=\frac{1}{|{\cal B}|^{2}}\sum_{b,b^{\prime}}\langle\eta_{ij}^{b}\eta_{kl}^{b^{\prime}}\rangle=\frac{1}{|{\cal B}|}\delta_{ik}\delta_{jl}.$ (3.35)
@@
$W_{ij}\to W_{ij}^{\prime}=W_{ij}-\alpha\frac{\delta{\cal L}}{\delta W_{ij}}+\frac{\alpha}{\sqrt{|{\cal B}|}}\sqrt{A_{ij}}\eta_{ij},\qquad\qquad\eta_{ij}\sim{\cal N}(0,1),$ (3.36)
which, as stated, is modelled explicitly to be consistent with the arguments given in Sec. 2.
We have applied this model to the case of $N=2$, with the teacher matrix diagonal with eigenvalues $\kappa_{1,2}=\kappa\pm\delta\kappa/2$, and $A_{ij}=A(1+\delta_{ij})$ (see App. A). The eigenvalues of $W$ then satisfy
$\lambda_{1}\to\lambda_{1}^{\prime}=\lambda_{1}-\alpha(\lambda_{1}-\kappa_{1})+\frac{\alpha^{2}}{|{\cal B}|}\frac{A}{\lambda_{1}-\lambda_{2}}+\frac{\alpha}{\sqrt{|{\cal B}|}}\sqrt{2A}\eta_{1},$ (3.37)
$\lambda_{2}\to\lambda_{2}^{\prime}=\lambda_{2}-\alpha(\lambda_{2}-\kappa_{2})+\frac{\alpha^{2}}{|{\cal B}|}\frac{A}{\lambda_{2}-\lambda_{1}}+\frac{\alpha}{\sqrt{|{\cal B}|}}\sqrt{2A}\eta_{2}.$ (3.38)
We have solved Eq. (3.36) numerically and varied the learning rate $\alpha$, mini-batch size $|{\cal B}|$ and noise strength $A$ independently. We have constructed the distribution for the level spacing and for the spectral density, and confirmed that in all cases these distributions can be fitted by the Wigner surmise and the Wigner semi-circle (for $N=2$).
@@
$\langle S\rangle-\delta\kappa-\frac{2\alpha A}{|{\cal B}|}\left\langle\frac{1}{S}\right\rangle=0.$ (3.39)
@@
$\langle S\rangle=\frac{1}{2}\left(\delta\kappa+\sqrt{\delta\kappa^{2}+\frac{4\pi\alpha A}{|{\cal B}|}}\right).$ (3.40)
@@
$\langle S\rangle=\sqrt{\frac{\pi\alpha A}{|{\cal B}|}},$ (3.41)
@@ 4 Summary and outlook
Moving forward, it is important to test these ideas in more complicated architectures, including neural networks and transformers. To be able to observe universal features, it may be required to use spectral unfolding, see e.g. Refs. [24, 25, 18]. The coupling between different weight matrices across layers as well as the use of adaptive learning rates, such as Adam [26], may also influence the manner in which universal features appear. Work along these directions is currently in progress.
@@
Research Data and Code Access – The code and data used for this manuscript will be made available in v2 of the arXiv version.

### v2

@@ III.2 Universal predictions
Next we turn to the spectral density (77) for each of the four doubly-degenerate modes. The results for two of these are shown in Fig. 5 (note that these are the second-to-lowest and second-to-highest histograms previously shown in Fig. 3). The lines are fits to Eq. (78) with the position $x_{m,i}$ and width $\sigma_{i}$ as free parameters. It is clear that the histograms are not described by simple Gaussians, but are broader and flatter. Non-Gaussianity can be checked by computing the Binder cumulant $U_{4}$ of the distribution, which equals (here $\delta x=x-x_{m}$)
$U_{4}\equiv\frac{\left\langle\delta x^{4}\right\rangle}{3\left\langle\delta x^{2}\right\rangle^{2}}-1=-\frac{4}{27}\approx-0.148$ (28)
for the Wigner semi-circle (with $N=2$), while it vanishes for a Gaussian distribution. For completeness, we note here that we have also fitted the distributions corresponding to a single, non-degenerate level (the ones furthest to the left and right in Fig. 3) and found these to be described by a Gaussian, as expected.
Importantly, from the general discussion in the previous section it follows that the higher the level of stochasticity (as indicated by $g_{i}^{2}$ and depending on learning rate and batch size), the wider the distributions are expected to be. We have hence repeated the exercise above for a range of learning rates, $0.01\leq\alpha\leq 0.1$, and mini-batch sizes, $4\leq|{\cal B}|\leq 64$. We will demonstrate below that the potentials $V_{i}(x_{i})$ have a global minimum at $x_{i}^{s}=\mu^{2}-\kappa_{i}$. The variance in the Coulomb gas $\sigma_{i}^{2}$ is then proportional to the ratio of the learning rate and batch size, see Eq. (18), and both the mean level spacing $\langle S\rangle$ and the width of the spectral density are expected to scale with $\sqrt{\alpha/|{\cal B}|}$ times a model and data dependent function of $\mu^{2}$ and $\kappa_{i}$.
@@
$\langle S\rangle,\sqrt{\pi}\sigma=a_{\rm fit}\sqrt{\frac{\alpha}{|{\cal B}|}\kappa_{i}^{2}\Omega_{i}}.$ (29)
A derivation of the non-universal factor $\kappa_{i}^{2}\Omega_{i}=(\mu^{2}-\kappa_{i})$, using an analysis of the variance of the gradient of the loss function, can be found in App. C. A linear dependence of $\langle S\rangle^{2}$ and $\sigma^{2}$ on $\alpha/|{\cal B}|$ is observed, consistent with the derivation in Sec. II.
@@
III.3 Non-universal dynamics
Next we turn to non-universal, model-dependent features, which are contained in the drift $K_{i}(x_{i})$ and the potential $V_{i}(x_{i})$. In Refs. Decelle_2021 ; Aarts:2023uwt it was shown that in the instantaneous eigenbasis the singular values $\xi_{i}$ of $W$ are subject to the equation
$\frac{d}{dt}\xi_{i}=\sigma_{h}^{2}\left(\frac{1}{\kappa_{i}}-\frac{1}{\mu^{2}-\sigma_{h}^{2}\xi_{i}^{2}}\right)\xi_{i},$ (30)
@@ III.3 Non-universal dynamics
$\frac{d}{d\tau}x_{i}=\left(\frac{1}{\kappa_{i}}-\frac{1}{\mu^{2}-x_{i}}\right)x_{i}.$ (31)
However, it should be clear from the previous discussion that this equation is not relevant for the evolution of the eigenvalues $x_{i}$ of $X$ in the realistic case, as it lacks both the Coulomb repulsion and the stochasticity. Instead, we should consider – see Eq. (63) –
$\frac{d}{d\tau}x_{i}=K_{i}(x_{i})+\sum_{j\neq i}\frac{g_{i}^{2}}{x_{i}-x_{j}}+\sqrt{2}g_{i}\eta_{i},$ (32)
@@
$K_{i}(x_{i})=\left(\frac{1}{\kappa_{i}}-\frac{1}{\mu^{2}-x_{i}}\right)x_{i}.$ (33)
@@
$\partial_{\tau}P(x,\tau)=\partial_{x}\left(g^{2}\partial_{x}-K(x)\right)P(x,\tau),$ (34)
with the drift
$K(x)=\left(\frac{1}{\kappa}-\frac{1}{\mu^{2}-x}\right)x.$ (35)
@@
$P_{s}(x)$ $=\frac{1}{Z}e^{-V(x)/g^{2}}$
$=\frac{1}{Z}\exp\left[\frac{1}{g^{2}}\left(\frac{x^{2}}{2\kappa}+x+\mu^{2}\log\left(\mu^{2}-x\right)\right)\right].$ (37)
@@
To analyse the properties of the time-dependent FPE, we cast the dynamics as a quantum-mechanical bound state problem Damgaard:1987rr . We factor out the square root of the stationary distribution
$P(x,\tau)=\sqrt{P_{s}(x)}\psi(x,\tau),$ (39)
and analyse the time dependence of $\psi(x,\tau)$. Taking the time derivative and using Eq. (34) then yields
$\partial_{\tau}\psi(x,\tau)=\Big{(}g^{2}\partial_{x}^{2}-\frac{1}{4g^{2}}\left[\partial_{x}V(x)\right]^{2}$
$\qquad+\frac{1}{2}\left[\partial_{x}^{2}V(x)\right]\Big{)}\psi(x,\tau)\equiv-2H_{\rm FP}\psi(x,\tau),$ (40)
@@
$L^{\dagger}=-g\partial_{x}+\frac{1}{2g}\partial_{x}V(x),\qquad L=+g\partial_{x}+\frac{1}{2g}\partial_{x}V(x),$ (42)
@@
$H_{\rm FP}\psi_{n}(x)=E_{n}\psi_{n}(x).$ (43)
The ground state, $\psi_{0}(x)$, with vanishing energy $E_{0}=0$, is determined by $L\psi_{0}(x)=0$, which yields $\psi_{0}(x)=\sqrt{P_{s}(x)}$, as it should be, see Eqs. (37) and (39).
@@
$\psi(x,\tau)=\psi_{0}(x)+\sum_{n>0}c_{n}\psi_{n}(x)e^{-2E_{n}\tau},$ (44)
where the coefficients $c_{n}$ are determined by the initial distribution $P(x,0)$, using relation (39).
@@
$H_{\rm FP}=-\frac{g^{2}}{2}\partial_{x}^{2}+U(x),\quad U(x)=\frac{1}{g^{2}}\left[U_{0}(x)+g^{2}U_{1}(x)\right],$ (45)
@@
$U_{0}(x)=\frac{1}{8}\left[\partial_{x}V(x)\right]^{2},\qquad U_{1}(x)=-\frac{1}{4}\partial_{x}^{2}V(x),$ (46)
@@
$\partial_{x}U_{0}(x)\big{|}_{x=x_{s}}=0,\qquad\partial_{x}V(x)\big{|}_{x=x_{s}}=-K(x_{s})=0,$ (47)
@@
$\partial_{x}^{2}U_{0}(x)\big{|}_{x=x_{s}}=\frac{\Omega^{2}}{4},\qquad\Omega\equiv\partial_{x}^{2}V(x)\big{|}_{x=x_{s}}=\frac{\mu^{2}-\kappa}{\kappa^{2}}.$ (48)
@@
$H_{\rm FP}=-\frac{g^{2}}{2}\partial_{x}^{2}+\frac{1}{2g^{2}}\left(\frac{\Omega}{2}\right)^{2}\left(x-x_{s}\right)^{2}-\frac{1}{2}\left(\frac{\Omega}{2}\right),$ (49)
@@
$E_{n}=\frac{\Omega}{2}n,\qquad\quad n=0,1,2,\ldots,$ (50)
and convergence to the stationary limit is determined by the smallest non-zero eigenvalue, $2E_{1}=\Omega$, as observed before Aarts:2023uwt . Also shown in Fig. 8 is the groundstate wave function, $\psi_{0}(x)=\sqrt{P_{s}(x)}$. We have a chosen a rather large value of $g^{2}$ here, such that it has a substantial width. Indeed, in the harmonic approximation, $g^{2}$ does not appear in the spectrum but controls the width of the wave function, and hence the distribution $P(x,t)$. Less stochasticity results in a narrower distribution centred around $x=x_{s}$, as one would expect.
@@
III.4 Teacher-student model
@@
$\mathbf{y}_{t}=V\mathbf{x},\qquad\qquad\mathbf{y}_{s}=W\mathbf{x},$ (51)
@@
${\cal L}(W)=\frac{1}{2}\left\langle||\mathbf{y}_{t}-\mathbf{y}_{s}||^{2}\right\rangle,$ (52)
where the brackets denote averaging over the input data. In the limit of an infinite amount of input data, $\langle x_{i}x_{j}\rangle=\delta_{ij}$, and the dynamics is deterministic. Some stochasticity can be introduced by considering batch updates. The amount of stochasticity is, however, not sufficient for the purpose of this discussion, since it is unavoidably suppressed when $W\to V$ as the training converges. We therefore introduce additional randomness by hand, in a manner that is consistent with the discussion in Sec. II, see in particular Eq. (5). We consider
$W_{ij}\to W_{ij}^{\prime}=W_{ij}-\alpha\frac{\delta{\cal L}}{\delta W_{ij}}+\alpha\sqrt{A_{ij}}\eta_{ij}^{\cal B}.$ (53)
@@
$\langle\eta_{ij}^{\cal B}\eta_{kl}^{\cal B}\rangle=\frac{1}{|{\cal B}|^{2}}\sum_{b,b^{\prime}}\langle\eta_{ij}^{b}\eta_{kl}^{b^{\prime}}\rangle=\frac{1}{|{\cal B}|}\delta_{ik}\delta_{jl}.$ (55)
@@
$W_{ij}\to W_{ij}^{\prime}=W_{ij}-\alpha\frac{\delta{\cal L}}{\delta W_{ij}}+\frac{\alpha}{\sqrt{|{\cal B}|}}\sqrt{A_{ij}}\eta_{ij},$ (56)
with $\eta_{ij}\sim{\cal N}(0,1)$, which, as stated, is modelled explicitly to be consistent with the arguments given in Sec. II.
We have applied this model to the case of $N=2$, with the teacher matrix diagonal with eigenvalues $\kappa_{1,2}=\kappa\pm\delta\kappa/2$, and $A_{ij}=A(1+\delta_{ij})$ (see App. D). The eigenvalues of $W$ then satisfy
$\lambda_{1}\to\lambda_{1}^{\prime}=$ $\,\lambda_{1}-\alpha(\lambda_{1}-\kappa_{1})+\frac{\alpha^{2}}{|{\cal B}|}\frac{A}{\lambda_{1}-\lambda_{2}}$
$+\frac{\alpha}{\sqrt{|{\cal B}|}}\sqrt{2A}\eta_{1},$ (57)
$\lambda_{2}\to\lambda_{2}^{\prime}=$ $\,\lambda_{2}-\alpha(\lambda_{2}-\kappa_{2})+\frac{\alpha^{2}}{|{\cal B}|}\frac{A}{\lambda_{2}-\lambda_{1}}$
$+\frac{\alpha}{\sqrt{|{\cal B}|}}\sqrt{2A}\eta_{2}.$ (58)
We have solved Eq. (56) numerically and varied the learning rate $\alpha$, mini-batch size $|{\cal B}|$ and noise strength $A$ independently. We have constructed the distribution for the level spacing and for the spectral density, and confirmed that in all cases these distributions can be fitted by the Wigner surmise and the Wigner semi-circle (for $N=2$).
@@
$\langle S\rangle-\delta\kappa-\frac{2\alpha A}{|{\cal B}|}\left\langle\frac{1}{S}\right\rangle=0.$ (59)
@@
$\langle S\rangle=\frac{1}{2}\left(\delta\kappa+\sqrt{\delta\kappa^{2}+\frac{4\pi\alpha A}{|{\cal B}|}}\right).$ (60)
@@
$\langle S\rangle=\sqrt{\frac{\pi\alpha A}{|{\cal B}|}},$ (61)
@@ IV Summary and outlook
Our results are rather general and can be applied to any model in which weight matrices are updated with a stochastic optimisation algorithm. The choice of architecture and loss function determine the non-universal aspects in the Coulomb potential, as well as the level of stochasticity. In some cases the resulting stochasticity is so small that is hardly noticeable; this was the case in the linear teacher-student model and for that reason we added noise by hand. The Gaussian RBM is stochastic by itself, due to the need to sample, and hence the noise inherent in the model is strong enough to observe the effects without any further addition. Moving forward, it is important to test these ideas in more complicated architectures, including neural networks and transformers. To be able to observe universal features, it may be required to use spectral unfolding, see e.g. Refs. Verbaarschot:2000dy ; AbulMagd:2014 ; Levi:2023qwg . The coupling between different weight matrices across layers as well as the use of adaptive learning rates, such as Adam kingma:2017 , may also influence the manner in which universal features appear. Work along these directions is currently in progress.
@@
Research Data and Code Access – The code and data used for this manuscript are available from Ref. zenodo .